A buffered solution contains 0.25 M NH3 (Kb = 1.8 x 10-5) and 0.40 M NH4Cl. a. Calculate the pH of this solution. b. Calculate the pH of the solution that results when 0.10 mole of gaseous HCl is added to 1.0 L of the buffered solution from part a.
1
Expert's answer
2020-03-25T12:46:52-0400

"pOH=pK_b+log[\frac{[NH_4^+]}{[NH_4OH]} ]"

"pOH=-log(1.8×10^{-5})+log[\frac{0.4}{0.25} ]"

"pOH=-log(1.8)+5+log(1.6)"

"pOH=5+log(\frac{1.6}{1.8})"

"pOH=5+log(16)-log(18)"

"pOH=5+4log(2)-log(9\u00d72)"

"pOH=5+4log(2)-log(2)-2log(3)"

"pOH=5+3\u00d70.301-2\u00d70.477"

"pOH=4.949"

"pH=9.051"

(b)"pOH=pK_b+log[\frac{[NH_4^+]}{[NH_4OH]} ]"

"pOH=-log(1.8×10^{-5})+log[\frac{0.5}{0.15} ]"

"pOH=-log(1.8×10^{-5})+log[\frac{0.5}{0.15} ]"

"pOH=5-log(1.8)+log[\frac{10}{3} ]"

"pOH=5-log(1.8)+log[10 ]-log(3)"

"pOH=5-log(18)+2-log(3)"

"pOH=5-log(2)+2-3log(3)"

"pOH=5.27"

"pH=8.73"